# Functional Liftings of Restricted Geometric Inequalities

### v1

@@ 2. Generalized Sup-Convolutions
$\geq\mathcal{M}_{\alpha}^{(t)}(\|f^{-}_{i}\|_{p_{i}})+\mathcal{M}_{\alpha}^{(t)}(\|f^{+}_{i}\|_{p_{i}})$
$=\mathcal{M}_{\alpha}^{(t)}(\lambda\|f_{i}\|_{p_{i}})+\mathcal{M}_{\alpha}^{(t)}((1-\lambda)\|f_{i}\|_{p_{i}})$
$=\mathcal{M}_{\alpha}^{(t)}(\|f_{i}\|_{p_{i}}),$
as desired. ∎
Concerning operations in line with our concentrations are the articles [28, 29]. In [28] Gardner, Hug, and Weil conducted a systemic study of operations between sets in ${\mathbb{R}}^{n}$ with a particular emphasis on compact convex sets in ${\mathbb{R}}^{n}$. Among these operations are the Minkowski sum, and $L_{p}$-Minkowski sum (see Section 5), which we consider here. Additionally, the article [29] of Gardner and Kiderlin concerns what they called “point-wise, homogeneous, monotonic, and associative” between functions defined on ${\mathbb{R}}^{n}$, and they characterize such operations as representable by $40$ explicitly given types. In particular, [29, Section 11] concerns a characterization of the supremal convolution appearing in (6) on the class of logarithmically concave functions on ${\mathbb{R}}^{n}$, denoted $\text{LC}({\mathbb{R}}^{n})$. They show, [29, Theorem 11.1], that any operation $\star:\text{LC}({\mathbb{R}}^{n})\times\text{LC}({\mathbb{R}}^{n})\to\text{LC}({\mathbb{R}}^{n})$ verifying certain properties, in addition to the above ones, must arise as a sup-convolution. Though most of our focus in this work is in Euclidean space, the framework for our “generalized sup-convolution” is much more flexible, approaching inequalities in measure spaces, both geometric and functional. An analogous classification of the seems beyond reach currently.
@@ 3. Dimensional Functional Brunn-Minkowksi inequalities in measure spaces
In this section we will extend several Brunn-Minkowski type inequalities from recent literature to functional versions. To start, as mentioned in the introduction the following result was conjectured by Gardner and Zvavitch in [30] and confirmed by Eskenazis and Moschidis in [24].
Theorem A (Eskenazis & Moschidis, [24]).
Let $\gamma$ be the standard Gaussian probability measure on ${\mathbb{R}}^{n}$ with density given by $\frac{d\gamma}{dx}=(2\pi)^{-\frac{n}{2}}e^{-\frac{|x|^{2}}{2}}$, For $t\in(0,1)$ and $A,B$ origin symmetric convex bodies in $\mathbb{R}^{n}$, it holds that
$\gamma((1-t)A+tB)\geq\mathcal{M}^{(t)}_{\frac{!}{n}}(\gamma(A),\gamma(B)).$
We prove the following Borell-Brascamp-Lieb type inequality holds for the Gaussian measure (namely Theorem 1.1) when restricting ourselves to even unimodal functions; this constitutes an improvement of the Prékopa-Leindler inequality (3).
@@
is a sup-convolution. Thus, in view of Theorem 2.4, our result holds if it holds for scaled indicator functions. Suppose that $f=a\mathbbm{1}_{A}$ and $g=b\mathbbm{1}_{B}$ for $A$ and $B$ convex bodies. For the remainder of the proof, we will always assume that $f$ and $g$ are of the above form. Observe that
$\gamma(f\square g)$ $=\mathcal{M}_{-\frac{1}{n}}^{(t)}(a,b)\gamma((1-t)A+tB)$
$\geq\mathcal{M}_{-\frac{1}{n}}^{(t)}(a,b)\mathcal{M}_{\frac{1}{n}}^{(t)}(\gamma(A),\gamma(B))$
$\geq\min\{a\gamma(A),b\gamma(B)\},$
@@
$=\mathcal{M}_{\alpha}^{(t)}\left(\frac{f(x)}{\gamma(f)}\gamma(f),\frac{g(y)}{\gamma(g)}\gamma(g)\right)$
$\geq\mathcal{M}_{-\frac{1}{n}}^{(t)}\left(\frac{f(x)}{\gamma(f)},\frac{g(y)}{\gamma(g)}\right)\mathcal{M}_{r}^{(t)}(\gamma(f),\gamma(g)).$
Applying the $\alpha=-\frac{1}{n}$ result to this inequality completes the proof of the theorem. ∎
This can be extended by the same method, using the following result.
Theorem B (Cordero-Erausquin & Rotem, [20]).
Let $w$ be an increasing function such that $s\mapsto w(e^{s})$ is convex, and $\mu$ the measure on $\mathbb{R}^{n}$ with density $\frac{d\mu}{dx}=e^{-w(|x|)}$. Then for symmetric convex bodies $A$ and $B$
$\mu((1-t)A+tB)\geq\mathcal{M}_{\frac{1}{n}}^{(t)}(\mu(A),\mu(B)).$
Theorem 3.1.
For $t\in(0,1)$, $\alpha\geq-\frac{1}{n}$ and $\mu$ a measure on $\mathbb{R}^{n}$ given by a density $\frac{d\mu}{dx}=e^{-w(|x|)}$ for $w$ increasing with $s\mapsto w(e^{s})$ convex, then if
$h((1-t)x+ty)\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y))$
holds for $h,f,g:\mathbb{R}^{n}\to[0,\infty)$, with $f,g$ symmetric unimodal functions, we have
$\mu(h)\geq\mathcal{M}_{r}^{(t)}(\mu(f),\mu(g))$
for $r=\frac{\alpha}{1+n\alpha}$.
@@
The proof follows by the same arguments as that of Theorem 1.1 while applying Theorem B. ∎
@@
Let $V\colon{\mathbb{R}}^{n}\to[0,\infty)$ be $s$-homogeneous convex function for some $1<p<\infty$, and let $\nu$ be the measure with density $\frac{d\nu}{dx}=e^{-V}$, and let $t\in(0,1)$ and $\alpha\geq\frac{1-s}{sn}$. If $h,f,g:\mathbb{R}^{n}\to[0,\infty)$ with $f,g$ star-unimodal, are a triple of functions satisfying
@@
where $r=\frac{\alpha(s-1)}{\alpha sn+s-1}$.

### v3

@@ 2. Generalized Sup-Convolutions
$\geq\mathcal{M}_{\alpha}^{(t)}(\|f^{-}\|_{1})+\mathcal{M}_{\alpha}^{(t)}(\|f^{+}\|_{1})$
$=\mathcal{M}_{\alpha}^{(t)}(\lambda\|f\|_{1})+\mathcal{M}_{\alpha}^{(t)}((1-\lambda)\|f\|_{1})$
$=\mathcal{M}_{\alpha}^{(t)}(\|f\|_{1}),$
as desired. Let now $p_{i}>1$ for all $i\in I\subset\{1,\ldots,n\}$ and $q_{i}$ be the dual exponent of $p_{i},$ for $i\in I$. We may write
$\mathcal{M}_{\alpha}^{(t)}(\|f_{i}\|_{p_{i}})=\sup_{\|g_{i}\|_{q_{i}}=1,i\in I}\mathcal{M}_{\alpha}^{(t)}\left(\left(\int f_{i}g_{i}d\mu_{i}\right)_{i\in I},\left(\|f_{i}\|_{1}\right)_{i\in I^{c}}\right).$
Thus (15) is equivalent to the fact that for each $g_{i}\in L^{q_{i}}(\mu_{i})$, with $\|g_{i}\|_{q_{i}}=1$ it holds that
(16) $\nu(\square f)\geq\mathcal{M}_{\alpha}^{(t)}\left(\left(\int f_{i}g_{i}d\mu_{i}\right)_{i\in I},\left(\|f_{i}\|_{1}\right)_{i\in I^{c}}\right).$
For each $i\in I$, let $g_{i}\in L^{q_{i}}(\mu_{i})$, with $\|g_{i}\|_{q_{i}}=1$ and consider the measures $\tilde{\mu}_{i}:=g_{i}\mu_{i}$, when $i\in I$. Then (14) implies that
$\nu(\square a)\geq\mathcal{M}_{\alpha}^{(t)}\left(\left(\|a_{i}\|_{L^{1}(\tilde{\mu}_{i})}\right)_{i\in I},\left(\|a_{i}\|_{1}\right)_{i\in I^{c}}\right),$
for each $a\coloneqq(a_{1}\mathbbm{1}_{A_{1}},\dots,a_{n}\mathbbm{1}_{A_{n}})$ for $a_{i}\in(0,\infty)$ and $A_{i}\in\mathcal{A}_{i}$. Since for each $i\in I$, $f_{i}\in L^{1}(\tilde{\mu}_{i})$ and for $i\in I^{c}$, $f_{i}\in L^{1}(\mu_{i})$, the previous case where all $p_{i}$’s were equal to $1$ gives (16) and thus concludes the proof for $p\in[1,\infty)^{n}$.
Concerning operations in line with our concentrations are the articles [25, 26]. In [25] Gardner, Hug, and Weil conducted a systemic study of operations between sets in ${\mathbb{R}}^{n}$ with a particular emphasis on compact convex sets in ${\mathbb{R}}^{n}$. Among these operations are the Minkowski sum, and $L_{p}$-Minkowski sum (see Section 4), which we consider here. Additionally, the article [26] of Gardner and Kiderlin concerns what they called “point-wise, homogeneous, monotonic, and associative” between functions defined on ${\mathbb{R}}^{n}$, and they characterize such operations as representable by $40$ explicitly given types. In particular, [26, Section 11] concerns a characterization of the supremal convolution appearing in (9) on the class of logarithmically concave functions on ${\mathbb{R}}^{n}$, denoted $\text{LC}({\mathbb{R}}^{n})$. They show, [26, Theorem 11.1], that any operation $\star:\text{LC}({\mathbb{R}}^{n})\times\text{LC}({\mathbb{R}}^{n})\to\text{LC}({\mathbb{R}}^{n})$ verifying certain properties, in addition to the above ones, must arise as a sup-convolution. Though most of our focus in this work is in Euclidean space, the framework for our “generalized sup-convolution” is much more flexible, approaching inequalities in measure spaces, both geometric and functional. An analogous classification of the seems beyond reach currently.
@@ 3. Dimensional Functional Brunn-Minkowksi inequalities in measure spaces
In this section we use Theorem 2.4 to obtain functional versions of dimensional Brunn-Minkowski inequalities. We will start with the proof of Theorem 1.1, for which the main result of [17] is required.
Theorem A (Cordero-Erausquin & Rotem, [17]).
Let $w:[0,\infty)\to(-\infty,\infty]$ be an increasing function such that ${\mathbb{R}}^{n}\ni s\mapsto w(e^{s})$ is convex, and $\mu$ the measure on $\mathbb{R}^{n}$ with density $\frac{d\mu}{dx}=e^{-w(|x|)}$. Then for symmetric convex bodies $A$ and $B$
$\mu((1-t)A+tB)\geq\mathcal{M}_{\frac{1}{n}}^{(t)}(\mu(A),\mu(B)).$
@@
is a sup-convolution. Thus, in view of Theorem 2.4, our result holds if it holds for scaled indicator functions. Suppose that $f=a\mathbbm{1}_{A}$ and $g=b\mathbbm{1}_{B}$ for $A$ and $B$ symmetric convex bodies. For the remainder of the proof, we will always assume that $f$ and $g$ are of the above form. Observe that
$\mu(f\square g)$ $=\mathcal{M}_{-\frac{1}{n}}^{(t)}(a,b)\mu((1-t)A+tB)$
$\geq\mathcal{M}_{-\frac{1}{n}}^{(t)}(a,b)\mathcal{M}_{\frac{1}{n}}^{(t)}(\mu(A),\mu(B))$
$\geq\min\{a\mu(A),b\mu(B)\},$
@@
$\geq\mathcal{M}_{-\frac{1}{n}}^{(t)}\left(\frac{f(x)}{\mu(f)},\frac{g(y)}{\mu(g)}\right)\mathcal{M}_{r}^{(t)}(\mu(f),\mu(g)).$
Applying the $\alpha=-\frac{1}{n}$ result to this inequality completes the proof of the general case. For the converse we apply the functional inequality to the characteristic functions of two symmetric convex bodies. ∎
In [16] a different functional formulation of the dimensional Brunn–Minkowski inequality was given and reads as follows (with the notation used so far):
Theorem B (Cordero-Erausquin & Eskenazis, [16]).
Let $\mu$ be the rotationally invariant measure on ${\mathbb{R}}^{n}$ as previously. Let $\Omega\subset{\mathbb{R}}^{n+1}$ be convex, with the property that for each $t\in{\mathbb{R}}$, $\Omega_{t}:=\{x\in{\mathbb{R}}^{n}:(t,x)\in\Omega\}$ is symmetric, and $\Phi:\Omega\to[0,\infty)$ be concave, with $\Phi(t,\cdot)$ even for each $t\in{\mathbb{R}}$. Then
(17) $\phi(t):=\left(\int_{\Omega_{t}}\Phi(t,x)^{\frac{1}{\alpha}}\ d\mu(x)\right)^{\frac{\alpha}{1+n\alpha}},$
is concave on its support for each $\alpha>0$, whenever the integral converges.
Further, in [16] it was asked whether the function $\phi$ is convex, whenever $\alpha\in(-\frac{1}{n},0)$ and $\Phi$ is convex. Using Theorem 1.1 we can answer positively this question for the class of measures considered so far and offer another proof of [16, Theorem 1].
Corollary 3.1.
Let $\mu$ be the measure on $\mathbb{R}^{n}$, with density $e^{-w(|x|)}$, where $w:[0,\infty)\to(-\infty,\infty]$ is increasing and $\mathbb{R}\ni t\mapsto w(e^{t})$ is convex. Let $\Omega\subset{\mathbb{R}}^{n+1}$ be convex, with the property that for each $t\in{\mathbb{R}}$, $\Omega_{t}:=\{x\in{\mathbb{R}}^{n}:(t,x)\in\Omega\}$ is symmetric, and $\Phi:\Omega\to[0,\infty)$ a function with $\Phi(t,\cdot)$ even for each $t\in{\mathbb{R}}$. Consider the function
(18) $\phi(t):=\left(\int_{\Omega_{t}}\Phi(t,x)^{\frac{1}{\alpha}}\ d\mu(x)\right)^{\frac{\alpha}{1+n\alpha}}.$
  1. (1)
    If $\alpha>0$ and $\Phi$ is concave, then $\phi$ is concave on its support, whenever the integral converges.
  2. (2)
    If $\alpha\in(-\frac{1}{n},0)$ and $\Phi$ is convex, then $\phi$ is convex on its support, whenever the integral converges.
@@
Consider the functions:
$h(z):=\Phi(\lambda t+(1-\lambda)s,z)^{\frac{1}{\alpha}}\mathbbm{1}_{\Omega_{\lambda t+(1-\lambda)s}}(z),\ f(x):=\Phi(t,x)^{\frac{1}{\alpha}}\mathbbm{1}_{\Omega_{t}}(x)\text{ and }g(y):=\Phi(s,y)^{\frac{1}{\alpha}}\mathbbm{1}_{\Omega_{s}}(y).$
Then we have that
$h(\lambda x+(1-\lambda)y)\geq\mathcal{M}_{\alpha}^{(\lambda)}(f(x),g(y))$
and thus Theorem 1.1 gives that $\phi$ has the desired property in both cases. ∎
Aishwarya and Rotem [2] recently used a new approach for dimensional Brunn-Minkowski inequalities using tools from optimal transport and information theory. Their entropic method produced the following geometric result.
@@
An application of Theorem 2.4 for this class of sets, similar to the above proof, gives the following functional analogue.
@@
Let $V\colon{\mathbb{R}}^{n}\to[0,\infty)$ be $s$-homogeneous convex function for some $1<p<\infty$, and let $\nu$ be the measure with density $\frac{d\nu}{dx}=e^{-V}$, and let $t\in(0,1)$ and $\alpha\geq-\frac{p-1}{pn}$. If $h,f,g:\mathbb{R}^{n}\to[0,\infty)$ with $f,g$ star-unimodal, is a triple of functions satisfying
@@
where $r=\frac{\alpha(p-1)}{\alpha pn+p-1}$.